OpenStudy (anonymous):

help plz if interest is compounded annually how much is $1000 worth in 9 years at 9%. indicate the formula also

OpenStudy (anonymous):

A(9) = 1000(1+(0.09/1))^(1*9)

OpenStudy (anonymous):

A(9) = 1000(1.09)^9 A(9) = 1000*1.9

OpenStudy (anonymous):

A(9) = $1900

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

why do they never explain this formula?

OpenStudy (anonymous):

A(t) = P(1+(r/n))^(nt) is the formula

OpenStudy (anonymous):

now there we go! not bad! so: A=P (!+ r/100)^t correct @aballer101 ?

OpenStudy (anonymous):

thnks!

OpenStudy (anonymous):

all I want to do is to let them understand, I know you know it @aballer101

OpenStudy (anonymous):

you have probably studied geometric sequences, from which this derives! @baller101

OpenStudy (anonymous):

you could probably write a whole tutorial on it @baller101

OpenStudy (anonymous):

but if I didnt know, what is A?

OpenStudy (anonymous):

did you explain that?

OpenStudy (anonymous):

I know it is the principal after nine years, but the student doesn't know that

OpenStudy (anonymous):

we are here to helpo, so lets help them!

OpenStudy (anonymous):

explain what you mean

OpenStudy (anonymous):

as Albert Einstein said.. Make it simple but no simpler!

OpenStudy (anonymous):

n oww I want something more complex, mad, huh?

OpenStudy (anonymous):

jesus christ this seems like a lecture!OMG, my apologies

OpenStudy (anonymous):

Mr.Hyde has taken over!